OpenStudy (henrietepurina):

Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (henrietepurina):

f(g(x))= ((4x-3)+3)/4= 4x/4=x

OpenStudy (henrietepurina):

g(f(x))=4( (x+3)/4 )-3=x

OpenStudy (imstuck):

Do you have the individual functions without the compositions? Like what is f(x) and what is g(x)?

OpenStudy (henrietepurina):

what? im sotty i dont understand...

OpenStudy (henrietepurina):

@Hero

hero (hero):

I thought you finished this already.

OpenStudy (henrietepurina):

Yeah but desmos wont graph this :(

hero (hero):

I don't think you understand what you are doing. That's not what you were supposed to graph.

OpenStudy (henrietepurina):

what?

hero (hero):

You were supposed to graph g(x) and f(x)

OpenStudy (henrietepurina):

yeah i did... but it didnt recognise it :(

hero (hero):

Do you remember the expressions for f(x) and g(x)?

OpenStudy (henrietepurina):

yeah... f(g(x))= ((4x-3)+3)/4= 4x/4=x g(f(x))=4( (x+3)/4 )-3=x

OpenStudy (henrietepurina):

@hero ?

hero (hero):

Those are the "compositions" of f and g, not the original expressions of f(x) and g(x).

OpenStudy (henrietepurina):

so what do i do?

hero (hero):

Find the original expressions for f(x) and g(x). Go back to the previous question you posted to find them.

OpenStudy (henrietepurina):

can you help me do that?

OpenStudy (henrietepurina):

@radar

OpenStudy (henrietepurina):

can you help me bud?

OpenStudy (henrietepurina):

i really need help on this....

hero (hero):

Okay, I'll help you with the expressions for f(x) and g(x) although I am surprised you don't know what they are. We went over this in detail earlier.

OpenStudy (henrietepurina):

im very sorry...

hero (hero):

\[f(x) = \frac{x + 3}{4}\] \[g(x) = 4x - 3\]

hero (hero):

Do you remember those expressions?

OpenStudy (henrietepurina):

yeah...

OpenStudy (henrietepurina):

the g(x) can be put on a graph but what about f(x)

OpenStudy (henrietepurina):

ok i graphed them.... what does it mean by Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (henrietepurina):

@Hero @radar

hero (hero):

Basically, you have to start with x = 0 and then provide the corresponding y value. Then continue until you get to x = 5 providing the corresponding y values for each value of x.